Shortcut method to find out series in binomial theorem.

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago
(x+y)0 = 1
(x+y)1 = x + y
(x+y)2 = x2 + 2xy + y2
(x+y)3 = x3 + 3x2y + 3xy2 + y3
(x+y)4 = x4 + 4x3y + 6x2y2 + 4xy3 + y4
(x+y)5 = x5 + 5x4y + 10x3y2 +10x2y3 + 5xy4 + y5
1
            1   1
          1   2   1
        1   3   3   1
      1   4   6   4   1
    1   5  10  10   5   1
  1   6  15  20  15   6   1
1   7  21  35  35  21   7   1
using binomial triangle.